# SLAM&Render: A Benchmark for the Intersection Between Neural Rendering, Gaussian Splatting and SLAM https://samuel-cerezo.github.io/SLAM&Render

### v1

@@ II-A Datasets for SLAM
The EuRoC dataset [17] provides visual-inertial data from a MAV, comprising two batches: one collected in an industrial environment for evaluating visual SLAM algorithms, and the other in a room for 3D environment reconstruction. Several challenges such as varying image brightness and dynamic scenes, are included. The TUM VI dataset [19] is a visual-inertial dataset designed for evaluating odometry algorithms. It includes RGB images, photometric calibration, and synchronized IMU data. Ground-truth poses at the start and end of the sequences are provided. The dataset also includes evaluations of state-of-the-art methods. Depth images are not present in this collection. OpenLORIS-Scene dataset [20] consists of visual-inertial and odometry data to benchmark SLAM and scene understanding algorithms in everyday scenarios. For data collection a wheeled robot equipped with a Realsense D435i camera, encoders and a LiDAR sensor was used. In the context of synthetic datasets, the Replica dataset [21] contains 18 photorealistic indoor reconstructions at room and building scales. It features dense 3D meshes, HDR textures, and reflective surfaces. However, neither changes in illumination or rearrangements of objects within the same scene are included.
@@
Unlike the existing dataset, the one presented in this work offers a comprehensive and diverse set of sequences designed to benchmark methods in the intersection between SLAM and novel view rendering. The dataset includes 40 sequences with synchronized streams of RGB images, depth, IMU, robot kinematic data, and ground-truth poses. The sequences cover five distinct setups with both consumer and industrial objects, captured under four lighting conditions. Each scene includes separate trajectories for training and testing, and variations in object arrangement to evaluate generalization. To the best of our knowledge, this is the first dataset containing multi-sensor and kinematic data collected in sequences characterized by varying lighting conditions, object rearrangements, occlusions, and time-synchronization (see Table I).
@@ III-A Convention for rigid transformations
where $\bm{R}_{A}^{B}\in SO\left(3\right)$ and $\bm{p}_{A}^{B}\in\mathbb{R}^{3}$ describe the orientation and the position of the frame $A$ with respect to the reference frame $B$, respectively. Such rigid transformation will transform a 3D point $\bm{p}^{B}\in\mathbb{R}^{3}$ ($\tilde{\bm{p}}^{B}$ in homogeneous coordinates) to the reference frame $A$ as $\tilde{\bm{p}}^{A}=\bm{T}_{A}^{B}\tilde{\bm{p}}^{B}$
@@ IV The SLAM&Render Dataset
In the website we provide a collection of tools for data association, retrieval from ROS2 bag files, alignment, and synchronization.
@@ IV-A Data acquisition
All RGB-D data were recorded at the optimal depth resolution [^2] https://dev.intelrealsense.com/docs/tuning-depth-cameras-for-best-performance (848×480) and full frame rate (30 Hz) of the Intel RealSense D435i camera. The IMU data were acquired at a rate of 210 Hz, capturing synchronized readings from the gyroscope and accelerometer. All the dataset sequences were collected through a system based on Ubuntu 22.04 and ROS2 Humble. We used the software library Intel RealSense SDK 2.0333https://www.intelrealsense.com/sdk-2/ and its ROS2 wrapper to record the camera data. While the depth and color images are initially unaligned, as they are captured by two distinct sensors, we used the built-in feature of the Intel ROS2 wrapper to register the depth images to the color images. This is achieved by projecting the depth images to 3D and then back-projecting them into the color camera view. The joint angles of the robot manipulator were collected using ROS2 at a frequency of 25 Hz.
@@
The cold and warm lighting setups were obtained thanks to a Neewer 660 PRO LED lamp (50 W).
@@ IV-C Reference frames
The reference frames shown in Fig. 4 were used in the data collection process to align the data coming from the camera, robot, and MCS. These reference frames are denoted as follows:
@@ IV-D Motion Capture System (MCS)
Synchronization. The data from the MCS are initially unsynchronized. To ensure temporal alignment, for each sequence we retrieved the capture starting time provided by OptiTrack. Consequently, the ground-truth file we provide includes the corresponding Unix Epoch timestamp for each pose, calculated using using the capture starting time. Following the approach presented in [25], the synchronization was validated by comparing the angular velocity estimated from the MCS with the gyroscope measurements from the IMU using cross-correlation (see Fig. 5). The maximum at zero offset confirms correct alignment.
@@ IV-E Camera calibration
TABLE III: Extrinsic parameters (hand-eye transformation) of the camera, represented as a translation [$mm$] and quaternion.
@@
TABLE IV: Mean, standard deviation and standard error of calibration metrics. The last two rows show the results using all images with the estimated and factory intrinsic parameters.
To compare the accuracy of the calibration using both factory and estimated intrinsic parameters, the set of calibration images was divided into several batches (10 batches with 20 images in each), and the same calibration process was applied to each batch. The accuracy of the calibration was assessed using the metrics proposed in [27], which include translation error (Trans. [$mm$]), rotation error (Rot. [$deg.$]), and relative transformation error (Rel. transf. error [$a.u.$]).
@@ V-B Results
Novel View Synthesis. To highlight the usefulness of independent test camera trajectories, we evaluated the performance of two state-of-the-art baselines, Gaussian Splatting [8] and FeatSplat [30], this last one specifically designed to reduce viewpoint overfitting. In this experiment, setup-1, setup-3, and setup-4 were used under natural, warm, and dark lighting conditions. Baselines were initialized with random point-clouds and a black background, using 4 Spherical Harmonics bands for Gaussian Splatting, and 32-dimensional latent vectors and 1 hidden layer of 64 neurons MLP for FeatSplat. They were optimized for a total of 21k iterations, and we report metrics both at 7k and 21k iterations. The baselines were optimized on the respective train trajectories of each setup, skipping one every ten frames. We evaluate both with the common approach of using skipped frames from the train trajectory [4, 8], Table V, and using the independently recorded test trajectory of each setup, Table VI.
@@
The common evaluation indicates an improvement in rendering performance, for both baselines, through the complete optimization process, see Table V. Both baselines achieve their best performance under dark lighting conditions, which is probably due to using black as initial background color. Nevertheless, the test trajectory at SLAM&Render reveals that both baselines overfit to the training trajectory. Note how the rendering quality does not improve beyond the 7k iteration and even slightly declines in Table Table VI. Qualitative analysis in Fig. 7 also showcases the contrast between the high quality rendering of validation viewpoints and the artifacts generated in test viewpoints. This result validates the need for datasets with independent trajectories to evaluate novel view synthesis methods and shows that the classic approach of sampling evaluation frames from the training trajectory does not inform about trajectory overfitting.
@@
Camera tracking. To illustrate the advantage on using kinematic data, three approaches based on Gaussian Splatting SLAM [9] (MonoGS) were evaluated. In the first experiment, we used the default configuration for MonoGS. In the second experiment we incorporated the robot kinematic data directly as the camera motion. This approach is referred to as the Kinematics experiment. In the third experiment, robot kinematrics were incorporated as an initial seed for camera pose estimation within MonoGS. This is referred to as Kinem. seed experiment. Following the evaluation conducted by the authors of MonoGS in [9], the ATE RMSE for keyframes is reported in Table VII. For this experiment, setup-1, setup-3, and setup-4 were used under different light conditions while performing train and test trajectories. In all cases, it is evident that incorporating kinematic data improves camera tracking performance. However, the results also show that the naive approach of using robot kinematics as the initial seed of MonoGS decreases its accuracy. This result motivates the use of our SLAM&Render dataset for research on multimodal fusion within NeRF and Gaussian Splatting SLAM. Fig. 8 illustrates the camera tracking results for setup-4 under cold lighting condition. The trajectory using MonoGS is shown in Fig. 8a, while Fig. 8b shows the result using kinematic data is used as initial seed. Incorporating kinematic data allows drift correction in camera tracking.
Figure 8: Camera tracking result for setup-4 with cold light using a) MonoGS and b) MonoGS with kinematic data as initial seed. Dashed line represent the ground-truth while colored line represent the camera trajectory. Incorporating kinematic data allows drift correction in camera tracking.
@@ VI Conclusions
In this work we introduced SLAM&Render, a dataset for benchmarking methods that combine SLAM and novel view rendering. Our data includes 40 sequences of synchronized RGB and depth images, IMU readings, robot kinematic data, intrinsic and extrinsic camera parameters, and ground-truth pose data across diverse scenes and conditions. SLAM&Render addresses research challenges such as generalization across trajectories and robustness to scene and illumination changes. It aims to drive future research in learning-based and hybrid methods requiring accurate localization and multi-sensor data fusion, ultimately enhancing the deployment of systems that combine perception, mapping, and rendering.

### v2

@@ II-A Datasets for SLAM
The EuRoC dataset [17] provides visual-inertial data from a Micro Aerial Vehicle, comprising two batches: one collected in an industrial environment for evaluating visual SLAM algorithms, and the other in a room for 3D environment reconstruction. Several challenges such as varying image brightness and dynamic scenes, are included. The TUM VI dataset [19] is designed for evaluating odometry algorithms. It includes RGB images, photometric calibration, and synchronized IMU data. Ground-truth poses at the start and end of the sequences are provided. The dataset also includes evaluations of state-of-the-art methods. Depth images are not present in this collection. OpenLORIS-Scene dataset [20] consists of visual-inertial and odometry data to benchmark SLAM and scene understanding algorithms in everyday scenarios. For data collection a wheeled robot equipped with a Realsense D435i camera, encoders and a LiDAR sensor was used. In the context of synthetic datasets, the Replica dataset [21] contains 18 photorealistic indoor reconstructions at room and building scales. It features dense 3D meshes, HDR textures, and reflective surfaces. However, neither changes in illumination or rearrangements of objects within the same scene are included.
@@
Unlike the existing datasets, the one presented in this work offers a comprehensive and diverse set of sequences designed to benchmark methods in the intersection between SLAM and novel view rendering. The dataset includes 40 sequences with synchronized streams of RGB images, depth, IMU, robot kinematic data, and ground-truth poses. The sequences cover five distinct setups with both consumer and industrial objects, captured under four lighting conditions. Each scene includes separate trajectories for training and testing, and variations in object arrangement to evaluate generalization. To the best of our knowledge, this is the first dataset containing multi-sensor and kinematic data collected in sequences characterized by varying lighting conditions, object rearrangements, occlusions, and time-synchronization (see Table I).
@@ III-A Convention for rigid transformations
where $\bm{R}_{A}^{B}\in SO\left(3\right)$ and $\bm{p}_{A}^{B}\in\mathbb{R}^{3}$ describe the orientation and the position of the frame $A$ with respect to the reference frame $B$, respectively. Such rigid transformation will transform a 3D point $\bm{p}^{A}\in\mathbb{R}^{3}$ ($\tilde{\bm{p}}^{A}$ in homogeneous coordinates) to the reference frame $B$ as $\tilde{\bm{p}}^{B}=\bm{T}_{A}^{B}\tilde{\bm{p}}^{A}$
@@ IV The SLAM&Render Dataset
Our website offers a collection of tools for data association, retrieval from ROS2 bag files, alignment, and synchronization.
@@ IV-A Data acquisition
All RGB-D data were recorded at the optimal depth resolution [^2] https://dev.intelrealsense.com/docs/tuning-depth-cameras-for-best-performance (848×480) and full frame rate (30 Hz) of the Intel RealSense D435i camera. The IMU data were acquired at a rate of 210 Hz, capturing synchronized readings from the gyroscope and accelerometer. All the dataset sequences were collected through a system based on Ubuntu 22.04 and ROS2 Humble. We used the software library Intel RealSense SDK 2.0333https://www.intelrealsense.com/sdk-2/ and its ROS2 wrapper [^4] https://dev.intelrealsense.com/docs/ros2-wrapper to record the camera data. While the depth and color images are initially unaligned, as they are captured by two distinct sensors, we used the built-in feature of the Intel ROS2 wrapper to register the depth images to the color images. This is achieved by projecting the depth images to 3D and then back-projecting them into the color camera view. The joint angles of the robot manipulator were collected using ROS2 at a frequency of 25 Hz.
@@
[…]
@@ IV-C Reference frames
The reference frames shown in Fig. 4 were used in the data collection process to align the data coming from the camera, the robot and the MCS. These reference frames are described below.
@@ IV-D Motion Capture System (MCS)
Synchronization. The data from the MCS are initially unsynchronized. To ensure temporal alignment, for each sequence we retrieved the capture starting time provided by OptiTrack. Consequently, the ground-truth file we provide includes the corresponding Unix Epoch timestamp for each pose, calculated using the corresponding capture starting time. Following the approach presented in [25], the synchronization was validated by comparing the angular velocity estimated from the MCS with the gyroscope measurements from the IMU using cross-correlation (see Fig. 5). The maximum at zero offset confirms correct alignment.
@@ IV-E Camera calibration
TABLE III: Extrinsic parameters (hand-eye transformation) of the camera, represented as a translation vector in [$mm$] and a unit quaternion for rotation.
@@
TABLE IV: Mean, standard deviation and standard error of the calibration metrics. The last two rows show the results using all images with the estimated and factory intrinsic parameters.
To compare the accuracy of the calibration using both factory and estimated intrinsic parameters, the set of calibration images was divided into several batches (10 batches with 20 images in each), and the same calibration process was applied to each batch. The accuracy of the calibration was assessed using the metrics proposed in [27], which include translation error in milimeters (Trans. [$mm$]), rotation error in degrees (Rot. [$deg.$]), and unitless relative transformation error (Rel. transf. error [$a.u.$]).
@@ V-B Results
Novel View Synthesis. To highlight the usefulness of independent test camera trajectories, we evaluated the performance of two state-of-the-art baselines: Gaussian Splatting [8] and FeatSplat [30] (this last one specifically designed to reduce viewpoint overfitting). In this experiment, setup-1, setup-3, and setup-4 were used under natural, warm, and dark lighting conditions. Baselines were initialized with random point-clouds and a black background, using 4 Spherical Harmonics bands for Gaussian Splatting, and 32-dimensional latent vectors and 1 hidden layer of 64 neurons MLP for FeatSplat. They were optimized for a total of 21k iterations, and we report metrics both at 7k and 21k iterations. The baselines were optimized on the respective train trajectories of each setup, skipping one every ten frames. We evaluate both with the common approach of using skipped frames from the train trajectory [4, 8] (see Table V) and using the independently recorded test trajectory of each setup (see Table VI).
@@
The common evaluation indicates an improvement in rendering performance, for both baselines, through the complete optimization process (see Table V). Both baselines achieve their best performance under dark lighting conditions, which is probably due to using black as initial background color. Nevertheless, the test trajectory at SLAM&Render reveals that both baselines overfit to the training trajectory. Note how the rendering quality does not improve beyond the 7k iteration and even slightly declines in Table Table VI. Qualitative analysis in Fig. 7 also showcases the contrast between the high quality rendering of validation viewpoints and the artifacts generated in test viewpoints. This result validates the need for datasets with independent trajectories to evaluate novel view synthesis methods and shows that the classic approach of sampling evaluation frames from the training trajectory does not inform about trajectory overfitting.
@@
Camera tracking. To illustrate the advantage on using kinematic data, three approaches based on Gaussian Splatting SLAM [9] (MonoGS) were evaluated. In the first experiment, we used the default configuration for MonoGS. In the second experiment we incorporated the robot kinematic data directly as the camera motion. This approach is referred to as the Kinematics experiment. In the third experiment, robot kinematrics were incorporated as an initial seed for camera pose estimation within MonoGS. This is referred to as Kinem. seed experiment. Following the evaluation conducted by the authors of MonoGS in [9], the ATE RMSE for keyframes is reported in Table VII. For this experiment, setup-1, setup-3, and setup-4 were used under different light conditions while performing train and test trajectories. In all cases, it is evident that incorporating kinematic data improves camera tracking performance. However, the results also show that the naive approach of using robot kinematics as the initial seed of MonoGS decreases its accuracy. This result motivates the use of our SLAM&Render dataset for research on multimodal fusion within NeRF and Gaussian Splatting SLAM. Fig. 8 illustrates the camera tracking results for setup-4 under cold lighting condition.
Figure 8: Camera tracking result for setup-4 under cold light using a) MonoGS and b) MonoGS with kinematic data as initial seed. Dashed line represent the ground-truth while colored line represent the camera trajectory. Incorporating kinematic data allows drift correction in camera tracking.
The trajectory using MonoGS is shown in Fig. 8a, while Fig. 8b shows the result using kinematic data as initial seed. Incorporating kinematic data allows drift correction in camera tracking.
@@ VI Conclusions
In this work we introduced SLAM&Render, a dataset for benchmarking methods that combine SLAM and novel view rendering. Our data includes 40 sequences of synchronized RGB and depth images, IMU readings, robot kinematic data, intrinsic and extrinsic camera parameters, and ground-truth pose data across diverse scenes and conditions. SLAM&Render addresses research challenges such as generalization across trajectories and robustness to scene and illumination changes. By releasing robot kinematic data, it also enables the assessment of novel SLAM strategies when applied to robot manipulators. It aims to drive future research in learning-based and hybrid methods requiring accurate localization and multi-sensor data fusion, ultimately enhancing the deployment of systems that combine perception, mapping, and rendering.